OpenStudy (anonymous):

Find the inverse of the function. f(x) = 3x - 2 How so?

OpenStudy (anonymous):

solve the equton for x.

OpenStudy (anonymous):

then sweitch the variable s x and y in the equation.

OpenStudy (anonymous):

should go something like this.

OpenStudy (anonymous):

(f(x)+2)/3=x; swtch the x and y variables to get (s+2)/3=y=f(x); sorry the s should be an x.